Earlier reports Convert Day of Year

The Hubble Space Telescope

Unlocking the Secrets of the Universe

Daily Report #5800

Period Covered:
08:00 pm May 30, 2012 - 07:59 pm May 31, 2012
(DOY 152/0000z - 152/2359z)

Flight Operations Summary

Significant Spacecraft Anomalies: (The following are preliminary reports
of potential non-nominal performance that will be investigated.)


19251-0 - Clear COS Event Flag 3 for prop 12810 visits 14-15@152/1938z


FGS GSAcq		8		8
FGS REAcq		7		7
OBAD with Maneuver	5		5

The COS FUV Special Recovery activities (proposal 12810, visits 12 - 13) successfully completed 
today without incident.  The FUV was ramped to the 1st HV plateau (-4916V/-4869V (154/151)) similar 
to visit 6; however this time the QE grid was on, no elevated counts or HV current transients were 
observed in the engineering telemetry during or after the ramp.  The STScI and MOSES2 Engineering 
Teams have examined the relevant telemetry and confirmed all voltages, currents and temperatures 
were as expected. The voltage and current monitors matched the signatures seen in SMOV.  The COS 
Science Team at STScI has examined the science data and DCE memory dumps and everything is nominal. 
 No HV current transients or unexpected elevated count rates were observed in the DCE memory dumps. 
The science data consisted of 2 sets of dark and wavecal exposures. The exposures were collected 
immediately after ramping to HVNom (154/151) and an hour later at the same HV level. Comparisons 
with SMOV data taken indicate comparable behavior when the aging of the detector and temporal 
fluctuations are considered.  The Engineering and Science Teams have given a "Go" to proceed with 
visit 14 and 15. Visits 14 and 15 are scheduled for Friday June 1 (DOY 153).  The COS event flag 3 
was cleared @152/19:36 via OR 19251  -jb  

Program Principal Investigator Program Title
Science Observations Scheduled
12457 Marc Postman, Space Telescope Science Institute Through a Lens, Darkly - New Constraints on the Fundamental Components of the Cosmos
12473 David Sing, University of Exeter An Optical Transmission Spectral Survey of hot-Jupiter Exoplanetary Atmospheres
12487 Xiaohui Fan, University of Arizona Probing Population III Star Formation in a z=7 Galaxy
12534 Harry Teplitz, California Institute of Technology The Panchromatic Hubble Ultra Deep Field: Ultraviolet Coverage
12576 Paul Kalas, University of California - Berkeley Orbit determination for Fomalhaut b and the origin of the debris belt halo
Calibration Observations Scheduled
12351 Knox Long, Space Telescope Science Institute IR Persistence 18
12688 Tiffany Borders, Space Telescope Science Institute Cycle 17: UVIS Bowtie Monitor
12689 Tiffany Borders, Space Telescope Science Institute WFC3 UVIS CCD Daily Monitor
12691 Vera Kozhurina-Platais, Space Telescope Science Institute WFC3/UVIS internal CTE monitor
12742 Justin Ely, Space Telescope Science Institute CCD Dark Monitor Part 2
12744 Justin Ely, Space Telescope Science Institute CCD Bias and Read Noise Monitor-Part 2
12784 Elena Sabbi, Space Telescope Science Institute - ESA Characterization of UVIS Traps Via Charge-Injected Biases
Other Observations Scheduled
12810 Thomas Wheeler, Space Telescope Science Institute COS/FUV Special Recovery